OpenStudy (anonymous):

Question 8 of 10 The function c(h) = 10 • 2h – 1 gives the number of customers in a store in hour h. What does the value 10 particularly represent in this situation? (Points : 3) The number of customers increases by 10 each hour. The store was open for 10 hours. The store had 10 customers in Hour 1. The number of customers increases by 10% each hour.

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

isnt it the number of hours ?

OpenStudy (anonymous):

@Nnesha

OpenStudy (danjs):

c(h) = 10 * 2h - 1

OpenStudy (anonymous):

right

OpenStudy (danjs):

test each option, increases by 10 each hour? use h = 1 and h=2 and see if c(h) increased by 10

OpenStudy (danjs):

c(1) = 10*2(1) - 1 = 19 c(2) = 10 *2(2) - 1 = 39 increased by 20 in 1 hour

OpenStudy (danjs):

c(3) = 10*2(3)-1 = 59

OpenStudy (anonymous):

confused

OpenStudy (danjs):

i tested choice A. It said the number of customers increases by 10 per hour. c(1) = 19 c(2) = 39 False, From hour 1 to hour 2, the number of customers increased by 20, not 10.

OpenStudy (anonymous):

oooh now i understand

OpenStudy (anonymous):

so are we trying to determine how much it has increased

OpenStudy (danjs):

Option 2, The store was open for 10 hours? It doesn't say that anywhere. False

OpenStudy (anonymous):

oh im sorry it increased by 20

OpenStudy (anonymous):

and right

OpenStudy (danjs):

Option 3) The store had 10 customers at hour 1? c(1) = 20 * 2(1) - 1 = 19 false, there were 19 customers at hour 1

OpenStudy (anonymous):

oh wow your right i think this equation alone just confused me i wasnt really looking at the answers

OpenStudy (danjs):

does it actually look like c(h) = 10*(2h-1) ? or

OpenStudy (anonymous):

the h is above the 2 ill draw it

OpenStudy (anonymous):